OpenStudy (anonymous):

A cylindrical boiler is to have volume of 1.340 cu. ft. The cost of the metal sheets to make boiler should be minimum.what should be its base diameter? a. 7.08ft b. 10.95ft c. 8.08.ft d. 11.95ft

OpenStudy (anonymous):

boiler means its open at the top?

OpenStudy (anonymous):

diameter of the base should be equal to the height

OpenStudy (anonymous):

"boiler" in this case just means cylinder

OpenStudy (phi):

The idea is the same as for the rectangular box problem. But you use slightly different equations. You need the formula for the surface area of a cylinder (with a top and bottom), and the formula for the volume of a cylinder. Google always works if you don't know these. \( A= 2 \pi r^2 + 2\pi r h\) \( V= \pi r^2 h \)

OpenStudy (dumbcow):

interesting to minimize surface area i get diameter should be about 1.19 but thats not one of the choices

OpenStudy (radar):

Every boiler I've ever seen has both a top and bottom ( a containment for water), used to create steam and heat a house. There is one near where I live, that is used to heat pallets.

OpenStudy (anonymous):

this is what i hate from my prof if the answer is not in the choices we Put E

OpenStudy (dumbcow):

hmm are you sure d) isn't off by a decimal ?

OpenStudy (anonymous):

11.95 yup

OpenStudy (anonymous):

if i imagine a boiler i feel the diameter is around 7ft up

OpenStudy (anonymous):

http://ce-review.com/category/mathematics_surveying_transportation/mathematics-surveying-and-trasportation-engineering/ my question is the same here but im a i.t student not a civil engineer

OpenStudy (phi):

In that case, 1.340 cu. ft. should read 1340 cu. ft. If we use this (new and improved) value, we get one of your choices.

OpenStudy (anonymous):

yah in my prof question 1.340 but in the site question its 1340 LOL

OpenStudy (phi):

Can you solve the problem? Start with \[ V= \pi r^2 h \] and solve for h in terms of V and r put this expression for h into \[ A= 2 \pi r^2 + 2\pi r h \] now take the derivative wrt r, set to 0 and solve for r double r to find d

OpenStudy (anonymous):

i cant solve this 1 honestly

OpenStudy (dumbcow):

ahh i had a feeling it had to be d) yeah i was using 1.34

OpenStudy (anonymous):

can i tell my prof the question is wrong? not accurate? what do you think?

OpenStudy (anonymous):

=)

OpenStudy (dumbcow):

no its just a typo, answer is clearly 11.95

OpenStudy (phi):

I would solve it using 1340, and say that otherwise none of the choices match

OpenStudy (phi):

The real point is learning how to solve these problems, and then showing people that you can do it.

OpenStudy (radar):

@phi, Would you put h in terms of R?

OpenStudy (radar):

\[V \over r ^{2}\pi\]

OpenStudy (phi):

\( h= \frac{V}{\pi r^2} \)

OpenStudy (radar):

Thanks.

OpenStudy (phi):

yes, where V is a known constant.

OpenStudy (radar):

In this case the volume is given, but, we are unsure of the accuracy. or veracity???

OpenStudy (phi):

The volume as given (1.34 cu. ft.) is inconsistent with any of the diameters given as a choice. Dumbcow found d= 1.19 ft. But if we use 1340 (matches the problem steph posted above) we can match with answer d

OpenStudy (radar):

O.K. HooRay for dumbcow!!

OpenStudy (phi):

Maybe the . in 1.340 should be a comma 1,340. Who knows?

OpenStudy (radar):

You are probably correct, a typo

OpenStudy (anonymous):

i will punish my prof on thursday dont worry =)

OpenStudy (anonymous):

Joke

OpenStudy (phi):

The question remains: Steph, can you solve this problem?

OpenStudy (anonymous):

waaa

OpenStudy (anonymous):

phi when my prof giving this test homework he always smile to all the student u know y? because he knows we dont know this =) but where working hard to find the answer..

OpenStudy (anonymous):

hey step are prof are the same when i check your question here i find that the question is wrong.1.340 = 1340 just tell the prof.

OpenStudy (anonymous):

but if they said its correct it doesnt matter.

OpenStudy (phi):

well, if you have \[ A= 2 \pi r^2 + 2\pi r \cdot \frac{1340}{\pi r^2} \] can you simplify and then find the derivative with respect to r?

OpenStudy (anonymous):

king answer it